Skip to main content
ins.: 2021-11-07 15:06; ( Age: 2 yrs )
By: Scientists
Category: STAR
STAR

STAR: 2021-11-06

The drillers moved from site 2 (TSM02) to site 3 (TSM03) today.

The drillers moved from site 2 (TSM02) to site 3 (TSM03) today. They were very fortunate that the rain forecast was wrong today and equipment was moved without any issues. Work will be resumed on Monday Nov 8th.